Given that a+b=c, a+c=d, and b=c+d, where d is a positive integer, what is the greatest value of a?
A)-3 B)2 C)-2 D)1 E)-1

Answer by greenestamps(13200)



Given:
(1) a+b = c
(2) a+c = d
(3) b = c+d
(4) d is a positive integer

Substituting (3) into (1),

a+c+d = c
a+d = 0
a = -d

Since d is a positive integer, a is a negative integer. The greatest value for a is when d has its minimum value of 1; then a is -1.

ANSWER: E) -1